A woman invested $15,000, part at 7% simple annual interest and part at 8% annual interest. If she receives $1,140 interest per year, how much did she invest at 7%?
Answer by mananth(16946)

Part I 7.00% per annum ------------- Amount invested =x
Part II 8.00% per annum ------------ Amount invested = y
15000
Interest----- 1140.00

Part I 7.00% per annum ---x
Part II 8.00% per annum ---y
Total investment
x + 1 y= 15000 -------------1
Interest on both investments
7.00% x + 8.00% y= 1140
Multiply by 100
7 x + 8 y= 114000.00 --------2
Multiply (1) by -7
we get
-7 x -7 y= -105000.00
Add this to (2)
0 x 1 y= 9000
divide by 1
y = 9000
Part I 7.00% $ 6000
Part II 8.00% $ 9000

CHECK
6000 --------- 7.00% ------- 420.00
9000 ------------- 8.00% ------- 720.00
Total -------------------- 1140.00

m.ananth@hotmail.ca